# JamendoMaxCaps: A Large Scale Music-caption Dataset with Imputed Metadata

### v1

@@ Abstract
We introduce JamendoMaxCaps, a large-scale music-caption dataset featuring over 200,000 freely licensed instrumental tracks from the renowned Jamendo platform. The dataset includes captions generated by a state-of-the-art captioning model, enhanced with imputed metadata. We also introduce a retrieval system that leverages both musical features and metadata to identify similar songs, which are then used to fill in missing metadata using a local large language model (LLLM). This approach allows us to provide a more comprehensive and informative dataset for researchers working on music-language understanding tasks. We validate this approach quantitatively with five different measurements. By making the JamendoMaxCaps dataset publicly available, we provide a high-quality resource to advance research in music-language understanding tasks such as music retrieval, multimodal representation learning, and generative music models.
@@ I Introduction
To mitigate these limitations, we introduce JamendoMaxCaps [^1] Dataset: https://huggingface.co/datasets/amaai-lab/JamendoMaxCaps  [^2] Code: https://github.com/AMAAI-Lab/JamendoMaxCaps, a large-scale music-caption dataset comprising over 200,000 freely-licensed instrumental tracks from the renowned Jamendo platform, a source of user-contributed music. We leverage a state-of-the-art music captioning model [4] to generate natural language descriptions for each track. To address the issue of incomplete metadata, we propose a multimodal metadata imputation approach that utilizes audio features as well a local large language model to expand the incomplete metadata. This allows us to populate fields such as genre, tempo, and variable tags like mood and instrumentation, providing a more comprehensive and informative dataset for researchers working on music-language understanding tasks.
@@
  • We present JamendoMaxCaps, a large-scale music-caption dataset featuring over 200,000 freely-licensed instrumental tracks from the Jamendo platform, along with generated captions.
@@ II Related Work
In this section, we discuss recent text-music datasets, metadata imputation research as well as music feature retrieval systems. MusicCaps [1], one of the most popular text-music datasets, contains human-annotated captions for a limited set of around 5,000 music samples. Despite the high-quality descriptions, its small scale is not sufficient for data-hungry deep learning models. MusicBench [20] further augments MusicCaps 11x and modifying the captions accordingly using an LLM. Still, it is limited to around 52k instances, each 10s in duration. The WavCaps dataset [19] expanded upon this concept by collecting over 400,000 audio-caption pairs from multiple online sources, employing heuristic filtering methods to ensure relevance. However, WaveCaps is more focused on generic audio rather than musical sounds. The recently released SongDescriber dataset [17], contains 1.1k human-written text captions of 706 songs. The authors intended this dataset to be used as an evaluation set for text-to-music systems. SunoCaps [5] contains 256 prompt-generated music samples, showing an alternative way creating text-music dataset. Finally, Noice2music [10] introduced the MuLaMCap dataset (MuLan-LaMDA Music Caption) which consists of 400k music-text pairs created by annotating AudioSet content with a data mining pipeline, but the dataset is not publicly available.
@@ III-A Caption Generation
To manage the computational demands of captioning over 200,000 tracks, we employed batched inference and leveraged 4-bit quantization to reduce model memory requirements, with minimal impact on caption quality. We set a maximum output length of 128 tokens per caption, ensuring that roughly 90% of the generated text falls within this range, thus avoiding overly long or unwieldy outputs. Each 30-second clip was processed using a standardized prompt:
@@
Our caption generation pipeline proves to be robust at scale, ultimately producing captions for more than 1.6 million 30-second audio segments, providing a reliable basis for subsequent metadata imputation and retrieval systems, as well as for broader music-language modeling efforts.
@@ IV-A JamendoMaxCaps Dataset
To create the dataset, we first collected audio tracks and associated metadata from the Jamendo music platform, a widely used online repository of Creative Commons-licensed music. We used their publicly accessible API to download instrumental songs that were uploaded to the platform between 2014 and 2024 (both years included), downloading the raw audio tracks and extracting the corresponding metadata fields. This process yielded a total of 202,603 audio tracks with partial metadata. We then cleaned the dataset by removing tracks with very short durations, resulting in a final set of 202,203 tracks with an average duration of 4 minutes and 20 seconds. A brief overview of some of the properties of the dataset is presented in Table I.
@@
Finally, we apply the metadata imputation approach discussed in the previous section. We observe that since all songs in our dataset are instrumental, ‘lang’ (language) and ‘gender’ (gender of singer) are always empty. We discard these for the evaluation of the imputation task. Considering only the rest of the metadata fields, 187,765 songs had partial metadata, while the remaining 14,438 tracks had complete metadata information. We focus on imputing the missing metadata fields for the songs with partial metadata. To create the final dataset, we concatenated the songs, their corresponding captions, and the (imputed where applicable) metadata information. A couple of examples of metadata imputation are given in Table II.
@@
Furthermore, we provide a summary of the distribution of some key metadata fields in the dataset to offer additional insight into its composition. This includes an analysis of the distribution of genres, and speed characteristics of the songs, which can help researchers and practitioners better understand the dataset and its potential applications. Figure 3 shows that the metadata imputation process adds more examples of previously underrepresented categories, such as “Techno” and “Jazz,” while also retaining the representation of major genres “Electronic” and “Rock”. This imputation effectively fills in the missing metadata fields, resulting in a more comprehensive and informative dataset representation. Moreover, the prevalence of “Electronic” and “Rock” genres in the dataset suggests the inclusion of more contemporary musical works, which aligns with the time period of our data collection efforts (2014-2024). In the case of speed, the imputation process enhances the granularity of the metadata by introducing additional categories, including “Fast”, “Slow”, “Very Fast”, and “Moderate”, which were not initially present significantly in the original data. While “Medium” remains the predominant category, the imputed metadata offers a more nuanced distribution, capturing a broader range of tempo variations. Noteworthy, in the original data, missing/nothing was the third most common category, but the imputation process successfully eliminated this gap, adding valuable information to the dataset. Figure 5 illustrates a significant enhancement in the diversity and coverage of vartags, providing more representation of musical characteristics. While the original data concentrated on a few prevalent tags such as “peaceful” and “energetic,” the imputed metadata introduces a wider range of descriptors, including “relaxing,” “futuristic,” and “dreamy”, which better encapsulate the emotional and atmospheric elements of the music.
@@ IV-B Retrieval System Settings
Since Qwen2-Audio [4] is already a published and established model, we do not attempt to evaluate its performance. Instead, we focus on the strength of the retrieval system and the metadata imputation approach.
@@ IV-C Objective Evaluation
To evaluate the accuracy of the metadata imputation, we first identify the examples that have valid entries in the original metadata for the ‘genres’, ‘speed’, and ‘vartags’ fields. We then generate imputed values for these fields using our approach and compare them to the original values for 5,000 examples per field. As the primary metric, we employ BERT-Score  [^7] https://arxiv.org/pdf/1904.09675 to measure the semantic similarity between the original and imputed values. Since the imputation can generate multiple predictions, this allows us to establish a semantic similarity score between the original and imputed data. Additionally, we utilize n-gram metrics such as BLEU1 to BLEU4 (B1, B2, B3, B4)  [^8] https://aclanthology.org/P02-1040.pdf to assess the exact overlap between the original and imputed values. Please note that the primary purpose of BLEU scores is to compare machine translations with ground truth, rather than evaluating semantic similarity.
@@ IV-D Subjective Evaluation
To subjectively verify the quality of the imputation, we conduct a listening study using the PsyToolkit platform [^9] https://www.psytoolkit.org/ [23]. Study participants are asked to listen to 10 randomly selected audio files and rate the associated imputed metadata on four criteria: 1) the extent to which the metadata could have been generated by a human, 2) the accuracy of the imputed genres, 3) the accuracy of the imputed speed, and 4) the accuracy of the imputed vartags. A 7-point Likert scale [16] is used for scoring, with 1 being the lowest and 7 being the highest. A total of 15 participants took part in the listening study.   [^10] Questions and samples: https://github.com/ijcnn4600/jamendomaxcaps.
The results of the listening study are presented in Table V. We achieve consistently high scores, with all metrics rated above 5.09 out of 7, indicating a strong overall performance. ‘speed’ and ‘vartags’ matching received particularly impressive scores of 5.82 and 5.59 respectively, with low standard deviations (0.43, 0.45) suggesting reliable and consistent performance across different musical pieces. The high score for human-like generation (5.32$\pm$0.31) indicates that our system successfully captures the characteristics that human annotators would typically provide. Even the relatively lower score for ‘genres’ matching (5.09$\pm$0.69) remains above the median point of the scale, demonstrating competent genre classification capabilities. These results suggest that our approach could serve as a reliable tool for automatic metadata generation maintaining high-quality standards.
@@ Acknowledgment
This work has received support from SUTD’s Kickstart Initiative under grant number SKI 2021_04_06. We acknowledge the use of ChatGPT for grammar refinement and paraphrasing.

### v2

@@ Abstract
We introduce JamendoMaxCaps, a large-scale music-caption dataset featuring over 362,000 freely licensed instrumental tracks from the renowned Jamendo platform. The dataset includes captions generated by a state-of-the-art captioning model, enhanced with imputed metadata. We also introduce a retrieval system that leverages both musical features and metadata to identify similar songs, which are then used to fill in missing metadata using a local large language model (LLLM). This approach allows us to provide a more comprehensive and informative dataset for researchers working on music-language understanding tasks. We validate this approach quantitatively with five different measurements. By making the JamendoMaxCaps dataset publicly available, we provide a high-quality resource to advance research in music-language understanding tasks such as music retrieval, multimodal representation learning, and generative music models.
@@ I Introduction
To mitigate these limitations, we introduce JamendoMaxCaps [^1] Dataset: https://huggingface.co/datasets/amaai-lab/JamendoMaxCaps  [^2] Code: https://github.com/AMAAI-Lab/JamendoMaxCaps, a large-scale music-caption dataset comprising over 362,000 freely licensed instrumental tracks from the renowned Jamendo platform, a source of user-contributed music. We leverage a state-of-the-art music captioning model [4] to generate natural language descriptions for each track. To address the issue of incomplete metadata, we propose a multimodal metadata imputation approach that utilizes audio features as well a local large language model to expand the incomplete metadata. This allows us to populate fields such as genre, tempo, and variable tags like mood and instrumentation, providing a more comprehensive and informative dataset for researchers working on music-language understanding tasks.
@@
  • We present JamendoMaxCaps, a large-scale music-caption dataset featuring over 362,000 freely-licensed instrumental tracks from the Jamendo platform, along with generated captions.
@@ II Related Work
In this section, we discuss recent text-music datasets, metadata imputation research as well as music feature retrieval systems. MusicCaps [1], one of the most popular text-music datasets, contains human-annotated captions for a limited set of around 5,000 music samples. Despite the high-quality descriptions, its small scale is not sufficient for data-hungry deep learning models. MusicBench [20] further augments MusicCaps 11x and modifying the captions accordingly using an LLM. Still, it is limited to around 52k instances, each 10s in duration. The WavCaps dataset [19] expanded upon this concept by collecting over 400,000 audio-caption pairs from multiple online sources, employing heuristic filtering methods to ensure relevance. However, WaveCaps is more focused on generic audio rather than musical sounds. The recently released SongDescriber dataset [17], contains 1.1k human-written text captions of 706 songs. The authors intended this dataset to be used as an evaluation set for text-to-music systems. SunoCaps [5] contains 256 prompt-generated music samples, showing an alternative way creating text-music dataset. Finally, Noise2music [10] introduced the MuLaMCap dataset (MuLan-LaMDA Music Caption) which consists of 400k music-text pairs created by annotating AudioSet content with a data mining pipeline, but the dataset is not publicly available.
@@ III-A Caption Generation
To manage the computational demands of captioning over 362,000 tracks, we employed batched inference and leveraged 4-bit quantization to reduce model memory requirements, with minimal impact on caption quality. We set a maximum output length of 128 tokens per caption, ensuring that roughly 90% of the generated text falls within this range, thus avoiding overly long or unwieldy outputs. Each 30-second clip was processed using a standardized prompt:
@@
Our caption generation pipeline proves to be robust at scale, ultimately producing captions for more than 3 million 30-second audio segments, providing a reliable basis for subsequent metadata imputation and retrieval systems, as well as for broader music-language modeling efforts.
@@ IV-A JamendoMaxCaps Dataset
To create the dataset, we first collected audio tracks and associated metadata from the Jamendo music platform, a widely used online repository of Creative Commons-licensed music. We used their publicly accessible API to download instrumental songs that were uploaded to the platform between 2008 and 2023 (both years included), downloading the raw audio tracks and extracting the corresponding metadata fields. This process yielded a total of 362,649 audio tracks with partial metadata. We then cleaned the dataset by removing tracks with very short durations, resulting in a final set of 362,238 tracks with an average duration of 4 minutes and 12 seconds. A brief overview of some of the properties of the dataset is presented in Table I.
@@
Finally, we apply the metadata imputation approach discussed in the previous section. We observe that since all songs in our dataset are instrumental, ‘lang’ (language) and ‘gender’ (gender of singer) are always empty. We discard these for the evaluation of the imputation task. Considering only the rest of the metadata fields, 261,935 songs had partial metadata, while the remaining 100,303 tracks had complete metadata information. We focus on imputing the missing metadata fields for the songs with partial metadata. To create the final dataset, we concatenated the songs, their corresponding captions, and the (imputed where applicable) metadata information. A couple of examples of metadata imputation are given in Table II.
@@
Furthermore, we provide a summary of the distribution of some key metadata fields in the dataset to offer additional insight into its composition. This includes an analysis of the distribution of genres, and speed characteristics of the songs, which can help researchers and practitioners better understand the dataset and its potential applications. Figure 3 shows that the metadata imputation process adds more examples of previously underrepresented categories, such as “Techno” and “Jazz,” while also retaining the representation of major genres “Electronic” and “Rock”. This imputation effectively fills in the missing metadata fields, resulting in a more comprehensive and informative dataset representation. Moreover, the prevalence of “Electronic” and “Rock” genres in the dataset suggests the inclusion of more contemporary musical works, which aligns with the time period of our data collection efforts (2008-2023). In the case of speed, the imputation process enhances the granularity of the metadata by introducing additional categories, including “Fast”, “Slow”, “Very Fast”, and “Moderate”, which were not initially present significantly in the original data. While “Medium” remains the predominant category, the imputed metadata offers a more nuanced distribution, capturing a broader range of tempo variations. Noteworthy, in the original data, missing/nothing was the third most common category, but the imputation process successfully eliminated this gap, adding valuable information to the dataset. Figure 5 illustrates a significant enhancement in the diversity and coverage of vartags, providing more representation of musical characteristics. While the original data concentrated on a few prevalent tags such as “peaceful” and “energetic,” the imputed metadata introduces a wider range of descriptors, including “relaxing,” “futuristic,” and “dreamy”, which better encapsulate the emotional and atmospheric elements of the music.
@@ IV-B Retrieval System Settings
Since Qwen2-Audio [4] is an established model, we do not evaluate its performance. Instead, we focus on the strength of the retrieval system and the metadata imputation approach.
@@ IV-C Objective Evaluation
To evaluate the accuracy of the metadata imputation, we first identify the examples that have valid entries in the original metadata for the ‘genres’, ‘speed’, and ‘vartags’ fields. We then generate imputed values for these fields using our approach and compare them to the original values for 5,000 examples per field. As the primary metric, we employ BERT-Score [27] to measure the semantic similarity between the original and imputed values. Since the imputation can generate multiple predictions, this allows us to establish a semantic similarity score between the original and imputed data. Additionally, we utilize n-gram metrics such as BLEU1 to BLEU4 (B1, B2, B3, B4) [22] to assess the exact overlap between the original and imputed values. Please note that the primary purpose of BLEU scores is to compare machine translations with ground truth, rather than evaluating semantic similarity.
@@ IV-D Subjective Evaluation
To subjectively verify the quality of the imputation, we conduct a listening study using the PsyToolkit platform [^7] https://www.psytoolkit.org/ [24]. Study participants are asked to listen to 10 randomly selected audio files and rate the associated imputed metadata on four criteria: 1) the extent to which the metadata could have been generated by a human, 2) the accuracy of the imputed genres, 3) the accuracy of the imputed speed, and 4) the accuracy of the imputed vartags. A 7-point Likert scale [16] is used for scoring, with 1 being the lowest and 7 being the highest. A total of 15 participants took part in the listening study.   [^8] Questions and samples: https://github.com/AMAAI-Lab/JamendoMaxCaps.
The results of the listening study are presented in Table V. We achieve consistently high scores, with all metrics rated above 5.09 out of 7, indicating a strong overall performance. ‘speed’ and ‘vartags’ matching received particularly impressive scores of 5.82 and 5.59 respectively, with low standard deviations (0.43, 0.45) suggesting reliable and consistent performance across different musical pieces. The high score for human-like generation (5.32$\pm$0.31) indicates that our system successfully captures the characteristics that human annotators would typically provide. Even the relatively lower score for ‘genres’ matching (5.09$\pm$0.69) remains above the median point of the scale, demonstrating competent genre classification capabilities. These results suggest that our approach could serve as a reliable tool for automatic metadata generation.
@@ Acknowledgment
This work has received support from SUTD’s Kickstart Initiative under grant number SKI 2021_04_06 and MOE under grant number MOE-T2EP20124-0014. We acknowledge the use of ChatGPT for grammar refinement and paraphrasing.